ADNAN MUHAMMED

Grade 12,

Successive discount of A percent and B percent are equivalent to a single discount of ?

 
Successive discount of A percent and B percent are equivalent to a single discount of ?
 

Grade:8

3 Answers

Arun
5 years ago
Dear student
 
It will be
 
A + B - AB /100
 
Hope it helps
In case of any fofficudif please feel.free to ask again
 
Regards
Arun (askIITians forume xpert)
Sayantan Garai
5 years ago
If M is the marked price then the selling price is
 (1-A/100)(1-B/100)×M.
Thus if M=100,
SP=(1-A/100)(1-B/100)×100.
Hence, Discount=(M-SP)=100-SP
             => D(%)= (M-SP)/M ×100= 100-SP.
AKASH
4 years ago
Let suppose be the price of material to be100 
 For A%discount=100-100*A%
                             =100-A
for AGAIN B%discount=(100-A)-(100-A)*B%
                                        =100-A-B+(AB)/100
        Net discount=(INITIAL COST-FINAL
 
COST)*100/INITIAL COST
          = (100-100+A+B-(AB)/100)*100/100
           Finally=A+B-(AB/100)